OpenStudy (anonymous):

the length of the hypotenuse of a right triangle is 15 ft.the length if 9 ft

OpenStudy (anonymous):

In a right angled triangle, we can use pythagoras thoerem..

OpenStudy (anonymous):

\[(Hypotenuse)^2 = (perpendicular)^2 + (base)^2\]

OpenStudy (anonymous):

Hypotenuse is given as 15 ft and you can take other length ie 9 ft as perpendicular or base, it is on you..

OpenStudy (igreen):

(Perpendicular and Base are pretty much the same thing) :P

OpenStudy (anonymous):

\[15^2 = 9^2 + (base)^2\]

OpenStudy (igreen):

As in, it doesn't matter which order you put them.

OpenStudy (anonymous):

I have taken 9 ft as perpendicular there, you can take base as well..

OpenStudy (anonymous):

im not good at math I even have my mom helping me and we still cant get it ;-;

OpenStudy (igreen):

\(a^2 + b^2 = c^2\) Plug in what we know: \(a^2 + 9^2 = 15^2\) Just simplify 9^2 and 15^2.

OpenStudy (anonymous):

Also, we are not.. We are just trying to help you..

OpenStudy (igreen):

Can you simplify \(9^2\) and \(15^2\)? @kaikoa

OpenStudy (perl):

we can also call them 'legs' , and hypotenuse

OpenStudy (igreen):

\(9^2 \rightarrow 9 \times 9\) \(15^2 \rightarrow 15 \times 15\)

OpenStudy (anonymous):

I would use first side and other side.. :P

OpenStudy (anonymous):

goodmorning and I cant help you my math is terrible aloha

OpenStudy (anonymous):

you are helping yourself I think.. :P

OpenStudy (igreen):

So bad that you can't multiply \(9 \times 9\)?

OpenStudy (anonymous):

I can but since its 4 in the morning IM LAZY AT THIS POINT

OpenStudy (igreen):

If you can't multiply \(9 \times 9\), you're really hopeless.

OpenStudy (anonymous):

(Facedesk)

OpenStudy (anonymous):

Use that calculator.

OpenStudy (anonymous):

Multiply 9 x 9 and 15 x 15.

OpenStudy (anonymous):

What do you get?

OpenStudy (anonymous):

Follow his username @kaikoa answer his questions.